OpenStudy (anonymous):

Find the variation constant and an equation of variation where y varies directly as x and y=8 when x=4

OpenStudy (anonymous):

variation constant is k .... so y/x=k k=2 but I dont understand the equation part ??

OpenStudy (anonymous):

y=kx; k=1/2

OpenStudy (anonymous):

so the variation constant is 1/2 not 2????

OpenStudy (anonymous):

sorry. y varies directly x i.e y is directly proportion to x. hence y = kx; => k = y /x where x = 4 and y =8 , => 8/4 = 2.

OpenStudy (lilg132):

it is 2

OpenStudy (lilg132):

all you have to do is y / x which is 8 / 4 = 2 = k

OpenStudy (anonymous):

just testing. your confidence.

OpenStudy (anonymous):

okay thanks guys, I get it now

OpenStudy (lilg132):

dont forget the medal or good answer :p